OpenStudy (anonymous):

2x1/-6x1/-5 please help!!

OpenStudy (anonymous):

the / is a division sign

OpenStudy (anonymous):

Can u pls type it using the equation tab below. It is difficult to understand it like this

OpenStudy (anonymous):

well h/o.. write ur answer as a fraction in simplest form. does this help? 2 X 1 X 1 --- --- -6 -5

OpenStudy (anonymous):

The answer is 1/15

OpenStudy (anonymous):

great! now how ab

OpenStudy (anonymous):

When you have such multiplication of fractions, follow the rule : (Here N means numerator and D means denomination) N1 x N2 x N3 N1 x N2 x N3 -- -- -- = ------------- D1 D2 D3 D1 X D2 X D3

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

Now using this formula try to solve yourself and I'll check if it is correct

OpenStudy (anonymous):

do u multiply the -7 with the N or D?

OpenStudy (anonymous):

in the 2nd problem?

OpenStudy (anonymous):

You can treat it as -7 -- 1 So you multiply -7 with the numerator group

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

soo 98 -- -54

OpenStudy (anonymous):

Correct!! But something more needs to be done

OpenStudy (anonymous):

simplify? idk how to do that with this fraction

OpenStudy (anonymous):

First, we hv to reduce it to lowest terms Second, make the denominator positive

OpenStudy (anonymous):

OK, Both are even nos and hence divisible by 2 So go ahead and divide both by 2

OpenStudy (anonymous):

Then multiply this by -1/-1 to make the denominator positive

OpenStudy (anonymous):

Done????

OpenStudy (anonymous):

49 -1 --- x ___ -27 -1

OpenStudy (anonymous):

soo that means the answer is this? -49 --- 27

OpenStudy (anonymous):

BINGO !!!! You got the hang of it !!

OpenStudy (anonymous):

yay! is that simplest form?

OpenStudy (anonymous):

Now I'll teach you how to reduce it to lower terms easily

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Yes it is because 49 is 7 x 7 and 27 is 3 x 3 x 3 So both numbers do not hv any common factor left

OpenStudy (anonymous):

okay thank u

OpenStudy (anonymous):

your question was 2 x -7 x -7 --- --- 9 -6

OpenStudy (anonymous):

We modify it to 2 x -7 x -7 --- --- --- 9 1 - 6

OpenStudy (anonymous):

Now we can easily see that 2 in the numerator of first fraction and 6 in the denominator of third fraction are both divisible by 2 and so we divide them by 2 to get 1 x -7 x -7 --- --- --- 9 1 -3

OpenStudy (anonymous):

Following me???

OpenStudy (anonymous):

ok now i have another ? is -3(3x-u)+(2u-4x) simplified to -9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now you can clearly see that numerators on top and denominators at bottom cannot be cross -cut as they do not have common factors so now multiply all numerators and all denominators and basically shift the - sign to the numerator

OpenStudy (anonymous):

is that the simplified expression to that expression? yes.thank u for explaining that to me.

OpenStudy (anonymous):

Yes, when u hv many fractions, you can use cross-cutting to cancel out common factors Now to yr second question..... -3(3x-u)+(2u-4x) = -9x + 3u + 2u - 4x = -9x - 4x + 3u + 2u = - 13x + 5u

OpenStudy (anonymous):

How did u arrive at -9 ??

OpenStudy (anonymous):

-3(-3x-u)+(2u-4x) -3(-3x)+1+(-4x) -3+1+-7

OpenStudy (anonymous):

i did it wrong

OpenStudy (anonymous):

but i was supposed to distribute

OpenStudy (anonymous):

Yes, when you have an expression like -3(3x-u), you have to first use the DISTRIBUTIVE LAW to open the bracket and then solve further

OpenStudy (anonymous):

wait..it said that wasn't correct.

OpenStudy (anonymous):

What's the ans u hv????

OpenStudy (anonymous):

i got 5x-1u

OpenStudy (anonymous):

1u ? :|

OpenStudy (anonymous):

-3(-3x-u) + (2u-4x) (9x-3u) + (2u-4x) -1u + 5x 5x + -1u 5x-1u

OpenStudy (anonymous):

Then u gave me the wrong question the first time around!!!! Go and check....

OpenStudy (anonymous):

i did! i'm sorry!! ahaha i put 3 and it was supposed to be -3 i think

OpenStudy (anonymous):

but am i correct (with the problem i just did)

OpenStudy (anonymous):

Well, one can always make mistakes while typing...... No, your answer is not correct. See my working below... -3(-3x-u) + (2u-4x) = 9x + 3u + 2u - 4x = 9x - 4x + 3u + 2u = 5x + 5u = 5 (x + u)

OpenStudy (anonymous):

"5x-1u " :\ why is there a 1 in front of the u ?

OpenStudy (anonymous):

you write "5-u" , no need for the 1 , it just makes it look messy , whenever I mark maths if I see anyone write coeffiecent of 1 then they lose marks

OpenStudy (anonymous):

But her answer is wrong anyway.....

OpenStudy (anonymous):

yeh , doesnt mean you cant correct notation

OpenStudy (anonymous):

agreed :)

OpenStudy (anonymous):

But if you deduct marks for this, I feel u r too harsh on the students. Just putting a circle and remarking shud do th job

OpenStudy (anonymous):

yeah but for my learning purposes i sometimes put the -1 so i don't lose track.

OpenStudy (anonymous):

but anyway..i see what u did harkirat ... idk what the crap i did! haha

OpenStudy (anonymous):

No sweat, we all learn by making mistakes and correcting them

OpenStudy (anonymous):

yeh, probably for like year4 kids in primary you wouldnt mark down , but if they are in high school and writting coefficents of 1 then there is something wrong , it makes you looks stupid putting 1u , because that shows the marker that you dont know that 1 times u =u

OpenStudy (anonymous):

@elecengineer R u located in the US??

OpenStudy (anonymous):

australia

OpenStudy (anonymous):

R u an Indian by any chance??

OpenStudy (anonymous):

no, not asian either

OpenStudy (anonymous):

but alot of the electrical engineering lecturers at my uni are

OpenStudy (anonymous):

im white lol

OpenStudy (anonymous):

Just asking, hope u didn't mind

OpenStudy (anonymous):

you shouldn't call students stupid...they mean the same thing. -1u and -u...just because it's marked differently doesnt mean you should mark down a grade..the answer is not wrong technically. its like your punishing your students because their teacher is a neat freak.

OpenStudy (anonymous):

I'm an Indian and I know teachers here (in India) r very harsh with such things and penalise even small things

OpenStudy (anonymous):

C'mon lauren, thats not nice

OpenStudy (anonymous):

i dnt care what they say ^

OpenStudy (anonymous):

but hey i'm just a first year college student. what do i know? but anywho the answer 5x+5u was correct!!! yayy!! i'm sorry. i'm just the type of person who says what i feel. i'm sorry if i offended you. just different opinions that is all.

OpenStudy (anonymous):

but seriously, if you are in high school you shouldnt need to write the 1s

OpenStudy (anonymous):

you are first year of uni , 19yrs old and you still write the 1s , you should have dropped that a long time ago lol

OpenStudy (anonymous):

C'mon friends, let us change the topic

OpenStudy (anonymous):

Do u need to ask anything else lauren??

OpenStudy (anonymous):

agreed. i don't like negativity. and yes...one sec please.

OpenStudy (anonymous):

im not negative , just stating the facts, and I would be very surprised if a uni tutor doesnt mark down because of bad notation , its one thing to get the right answer, but you also need to express yourself confidently

OpenStudy (anonymous):

if you click the link it makes it bigger.

OpenStudy (anonymous):

^ thats first year uni math? cant be

OpenStudy (anonymous):

what do you mean by "college"

OpenStudy (anonymous):

i'm in Math098...i'm not the best math student and have never really been good at math or liked it...but it is a lower level of math.

OpenStudy (anonymous):

but i do go to a university.

OpenStudy (anonymous):

k... thats like the stuff we do in year8 in high school in australia

OpenStudy (anonymous):

lauren, u do it and tell me the ans I'll confirm if it is ok This is a simple one, just b careful while using the Dist. LAW

OpenStudy (anonymous):

i got -11y+8

OpenStudy (anonymous):

wow :p

OpenStudy (anonymous):

Bingo !! Its correct , the working as follows -4 (3y - 2) + y = -12y + 8 + y = - 11y + 8

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

I should go fetch a really hard version of these from one of my high maths text books , these are far too easy :p

OpenStudy (anonymous):

something much more complex

OpenStudy (anonymous):

k u go do that. lol

OpenStudy (anonymous):

Ok lauren, give this a try 2(x+ 3y) -7(2y - 5z) + 3(2x - 3y +z)

OpenStudy (anonymous):

Focus and you'll do it right

OpenStudy (anonymous):

k h/o